40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13453     Entry time: Thu Nov 23 18:03:52 2017     In reply to: 13452
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: Oplev "noise budget" 

Here are a couple of preliminary plots of the noise from a 20minute stretch of data - the new curve is the orange one, labelled sensing, which is the spectrum of the PIT/YAW error signal from the HeNe beam single bounce off a single steering mirror onto the QPD, normalized to account for the difference in QPD sum. The peaky features that were absent in the dark noise are present here.

I am a bit confused about the total sum though - there is ~2.5mW of light incident on the PD, and the transimpedance gain is 10.7kohm. So I would expect 2.5e-3 mW * 0.4A/W * 10.7 kV/A ~ 10.7V over 4 quadrants. The ADC is 16 bit and has a range +/- 10V, so 10.7 V should be ~35,000 cts. But the observed QPD sum is ~14,000 counts. The reflected power was measured to be ~250uW, so ~10% of the total input power. Not sure if this is factored into the photodiode efficiency value of 0.4A/W. I guess there is some fraction of the QPD that doesn't generate any photocurrent (i.e. the grooves defining the quadrants), but is it reasonable that when the Oplev beam is well centered, ~50% of the power is not measured? I couldn't find any sneaky digital gains between the quadrant channels to the sum channel either... But in the Oplev setup, the QPD had ~250uW of power incident on it, and was reporting a sum of ~13,000 counts with a transimpedance gain of 100kohm, so at least the scaling seems to hold...

I guess we wan't to monitor this over a few days, see how stationary the noise profile is etc. I didn't look at the spectrum of the intensity noise during this time.

Quote:

I've setup a test setup on the ITMY Oplev table. Details + pics to follow, but for now, be aware that

Here are some pics of the setup: https://photos.app.goo.gl/DHMINAV7aVgayYcf1.None of the existing Oplev input/output steering optics were touched. Steve can make modifications as necessary, perhaps we can make similar mods to the SRM Oplev QPD and the BS one to run the HeNe test for a few days...

 

 

Attachment 1: ITMY_P_noise.pdf  251 kB  | Hide | Hide all
ITMY_P_noise.pdf
Attachment 2: ITMY_Y_noise.pdf  250 kB  | Hide | Hide all
ITMY_Y_noise.pdf
ELOG V3.1.3-